Simplify and express the result in power notation with positive exponent:
 (1/2^3 )^2  

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1^2(2^3)^2
(a/b)^m = a^m/b^m
= 1/2^3×2= 1/2^6
[ (a^m)^n = a^mn ]

Thanks